a body of mass 2 kg moves vertically downwards with an acceleration 19.6 the force acting on the body simultaneously with the force of gravityis 9.8

Grade:11

2 Answers

Vikas TU
4 years ago
Force acting on the body will be equal to 
F + mg = ma
Insert the above value in the equation , 
We get F = 19.6 N 
Hope this helps 
Cheers
Khimraj
4 years ago
F = additional force acting on the body in down direction
m = mass of the body = 2 kg 
g = acceleration due to gravity = 9.8 ms⁻²
a = acceleration in down direction = 19.6 ms⁻²
force equation for the motion of stone in down direction can be given using newton's second law as
F + mg = ma 
inserting the above values
F + (2) (9.8) = (2) (19.6)
F = 19.6 N